Wes is mixing peanuts and chocolate pieces. Peanuts sell for $4.50 a pound and the chocolate sells for $6.50 a pound. How many pounds of chocolate mixes with 5 pounds to obtain a mixture that sells for $5.25 a pound?
Answer by alicealc(293)

chocolate = x pound

4.5*5 + 6.5*x = 5.25*(5+x)
22.5 + 6.5x = 26.25 + 5.25x
6.5x - 5.25x = 26.25 - 22.5
1.25x = 3.75
x = 3.75/1.25 = 3

so we need 3 pounds of chocolate